OpenStudy (anonymous):

how do you simplify 2/3*3/2 under the radical?

OpenStudy (lgbasallote):

\(\huge \sqrt{\frac{2}{3} \times \frac{3}{2}}?\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

multiply out and get sqrt(1) ?

OpenStudy (anonymous):

you can say that (2/3)^1/2 * (3/2)^1/2 when you put them side by side the 1st 3 in the denominator will cancell out with the 2nd 3 in the denominator and so will the 2's leaving you with 1^(1/2) which is 1